15.00 gm of a compound Na2SO4. xH2O  contains 2.3* 10 ^23 H2O molecules. Find the value of x.
How to solve this??

Grade:

1 Answers

Arun
5 years ago
1 mole of molecules contain 6.022x10^23 molecules (Avogadro's number)
 
Moles of H2O = 2.36× 10^23 / 6.022x10^23
                          = 0.392 moles

Molar mass of H2O = 18

Mass of H2O = 0.392x18 = 7.05g

Mass of Na2SO4 = 15 – 7.05 = 7.95g

Molar mass of Na2SO4 = 23x2 + 32 + 16x4 = 142

Moles of Na2SO4 = 7.95/142 = 0.0556

Mole ratio:
Na2SO4 = 0.0556/0.0556 = 1
H2O       = 0.392/0.0556 = 7

The value of x = 7

The formula of the compound is: Na2SO4 .7 H2O